OpenStudy (ellamoyseyuk):

You are making the kite shown at the right from five pairs of congruent panels. In parts (a)–(d) below, use the given information to find the side lengths of the kite’s panels. ABCD is a kite. EB=15 in., BC=25 in. The extended ratio XY:YZ:ZC is 3:1:4. EX is perpendicular BC, EX is parallel to YF, YF is parallel to GZ. a) triangle BEX, b) XEFY, c) YFGZ, d) triangle ZGC

OpenStudy (ellamoyseyuk):

EC^2 = 625 - 225 EC^2 = 400 EC = sqrt(400) EC = 20

OpenStudy (ellamoyseyuk):

EX^2 = 25^2 -20^2 EX^2= 625- 400 EX^2= 225 EX=sqrt (225) EX= 15

OpenStudy (ellamoyseyuk):

XC^2= 25^2-15^2 XC^2= 625-225 XC^2= 400 XC= sprt(400) XC= 20

OpenStudy (ellamoyseyuk):

How would you get XY:YZ:ZC and a) triangle BEX, b) XEFY, c) YFGZ, d) triangle ZGC

OpenStudy (phi):

I'm checking your work for EX. Finding EX is a bit tricky if we draw the shape this way